The equilibrium constant for the reaction HNO2 (aq) + H2 O ( ℓ) ↔ NO2¯ (g) + H3O+ (aq) is 4.3 x 10 − 4 at 25º C. Will nitrous acid spontaneously dissociate when:

a) [HNO2] = [NO2¯] = [H3O + ] = 1.0 M ?

b) [HNO2] = 1.0 M and [NO2¯ ] = [H3O + ] = 1.0 x 10− 5 M ?

Please help solve this question and include the process as well. Thanks.
Answer #1


Let's first understand that the spontaneity of the equilibrium reaction depends on equilibrium constant.
A spontaneous reaction is a reaction that will proceed without any outside energy or driving force. A spontaneous reaction has an equilibrium constant greater than 1. A reaction will be nonspontaneous if the equilibrium constant is less than 1.
The equilibrium constant is the value Kc when the reaction is at equilibrium. If the chemicals in the reaction are not at equilibrium, then the value obtained by the equilibrium law is called the reaction quotient. Q has the same form as the equilibrium law, except Kc is replaced by Q. Following are the properties derived from the definition of the Q:

If Q = Kc, the reaction is at equilibrium.
If Q does not change with respect to time, the reaction is at equilibrium and thus, Q =Kc.
If Q < Kc, the reaction will move to the right (the forward direction) in order to reach equilibrium.
If Q > Kc, the reaction will move to the left (the reverse direction) in order to reach equilibrium.

a) Q = ([NO2-][H3O+]) / ([HNO2]) = (1/1) = 1 so Q > Kc. Thus it is not spontaneous or it will favor backword reaction.

b) Q = ((1*10^-5)^2) / (1) = 1*10^-10, so Q < Kc. Thus the reaction will favor in forward direction to sponteneously dissociate the nitrous acid.